OpenStudy (cutestgirlaround):

In 1992, Jason bought a gallon of gas for $1.19. Yesterday, he bought a gallon of gas for $2.16. What is the percentage increase of the price of a gallon of gas from 1992 to yesterday? If necessary, round to the nearest tenth of a percent. 55.1% 18.5% 44.9% 81.5%

OpenStudy (decarr432):

What do you think the answer is?

OpenStudy (campbell_st):

use % increase = (change in price)/(initial price0 x 100

OpenStudy (cutestgirlaround):

c

OpenStudy (cutestgirlaround):

am i right

OpenStudy (decarr432):

Why do you think this is correct?

OpenStudy (cutestgirlaround):

c

OpenStudy (campbell_st):

what did you get for the change in price...?

OpenStudy (campbell_st):

the initial price is $1.19

OpenStudy (cutestgirlaround):

am i right @Decarr432

OpenStudy (campbell_st):

its quicker if you do the calculation.

OpenStudy (decarr432):

Why do you think this answer is the correct, please put forth effort :D

OpenStudy (campbell_st):

remember, you are being asked to compare it to the 1992 price... not yesterday's price

OpenStudy (campbell_st):

change = 2.16 - 1.19 initial price = 1.19 \[Change = \frac{change}{initial~price} \times 100\]

OpenStudy (katie_rogers101):

using this formula you can come to the conclusion that D is the answer. But keep in mind you need to learn and understand this formula incase you have testing on it.